# Examples of topologically unknotted tori

### v2

@@ 1.2. Seifert surface union a ribbon disc
is surjective. In particular, every ribbon disc is homotopy ribbon; see Gordon [Gor81]. Let $K$ be a knot in $S^{3}$ with Seifert surface $F$ and homotopy ribbon disc $D$. Then $\Sigma:=F\cup D\subseteq S^{4}$ is a $\mathbb{Z}$-surface. A slice derivative for a genus $g$ Seifert surface $F$ for a knot $K$ in $S^{3}$ is a collection of $g$ pairwise disjoint, homologically linearly independent simple closed curves that form a smoothly slice link and have self-linking zero [CHL10]. If $F$ has a slice derivative, we can perform surgery on $F$ in $D^{4}$ along a slice derivative and obtain a smooth slice disc $D$ for $K$. Then $F\cup D$ is smoothly unknotted; see Lemma 6.1. If $D$ is a ribbon disc that is not obtained from a Seifert surface $F$ by surgering along a slice derivative, Gabai has asked whether $F\cup D$ is topologically or smoothly unknotted. Kauffman conjectured that every genus one Seifert surface for a slice knot admits a slice derivative. Cochran and Davis [CD15] have shown that there exists a ribbon knot admitting a genus one Seifert surface $F$ that has no slice derivative. Moreover, there is an example where $F$ is the unique minimal genus one Seifert surface, up to isotopy. The union of $F$ and a ribbon disc is therefore a potentially smoothly knotted $\mathbb{Z}$-torus in $S^{4}$. We show that it is topologically unknotted.
@@ 1.3. Tori with four critical points
Let $\Sigma\subseteq S^{4}$ be a smoothly embedded torus with four critical points. Suppose that the single-variable Alexander polynomial $\Delta_{L}$ of the 2-component link $L:=\Sigma\cap S^{3}$ is zero. Then $\Sigma$ is topologically unknotted.
@@
Let $\Sigma\subseteq S^{4}$ be a torus with four critical points, and suppose that the link $L:=\Sigma\cap S^{3}$ is split. Then $\Sigma$ is smoothly unknotted.
@@ Acknowledgements
We would like to thank Anthony Conway, Chris Davis, David Gabai, Jason Joseph, Maggie Miller, András Stipsicz, Zoltán Szabó, and Ian Zemke for helpful discussions and suggestions. We are also particularly grateful to Brendan Owens, who pointed out a mistake in our proof of an earlier version of 1.4. MP was partially supported by EPSRC New Investigator grant EP/T028335/2 and EPSRC New Horizons grant EP/V04821X/2.
@@ 2. Standard intersection form implies unknotted
The elements of the standard basis $\{e_{i}\}_{i=1}^{2g}$ of $H_{2}(E_{\Sigma};\Lambda)\cong\Lambda^{2g}$ are represented by generically immersed spheres, which we can see as follows. Apply successively the Hurewicz theorem to the universal cover of $E_{\Sigma}$ to see that $\pi_{2}(E_{\Sigma})\cong H_{2}(E_{\Sigma};\Lambda)$, the density of smooth immersions in $C^{0}(S^{2},M)$ [Hir76, Theorems 2.2.6 and 2.2.12], and finally the density of generic immersions in the space of smooth immersions [GG73, Chapter III, Corollary 3.3]. Since the $\{e_{i}\}_{i=1}^{2g}$ are represented by generically immersed spheres, so are $\{e_{i}\otimes 1\}_{i=1}^{2g}\subseteq H_{2}(E_{\Sigma};\Lambda)\otimes_{\Lambda}\mathbb{Z}\cong\Lambda^{2g}\otimes_{\Lambda}\mathbb{Z}$.
@@
for some $a_{ij}\in\mathbb{Z}$ such that the matrix $(a_{ij})_{i,j=1}^{2g}$ is invertible over $\mathbb{Z}$. Hence, the elements $\{e_{i}^{\prime}:=\sum_{j}a_{ij}e_{j}\}_{i=1}^{2g}$ generate $H_{2}(E_{\Sigma};\Lambda)$, because $(a_{ij})_{i,j=1}^{2g}$ is also invertible over $\Lambda$. Furthermore, the $e_{i}^{\prime}$ map to the rim tori. By taking parallel copies of the $e_{j}$ and tubing together, each of the $e_{i}^{\prime}$ can also be represented by generically immersed spheres. ∎
@@
It will sometimes be helpful to compute the homological self-intersection numbers $\lambda([S_{i}],[S_{i}])\in\Lambda$ in Lemma 2.4 using the Wall self-intersection number $\mu(S_{i})$, so we recall this next; see [Wal99, Chapter 5]. Let $S\subseteq E_{\Sigma}$ be a $\mathbb{Z}$-trivial surface that is the image of a generic immersion into $E_{\Sigma}$; i.e. all self-intersections are transverse double points. For each double point $p\in S$, we obtain a sign $\varepsilon_{p}$ and a double point loop $\gamma_{p}$ in $S$, with the latter well-defined up to orientation and homotopy in $E_{\Sigma}$; see e.g. [${}^{+}$21, Section 11.3]. Thus each double point gives rise to a monomial ±$\pm t^{k}\in\Lambda/\langle t^{r}\sim t^{-r}:r\in\mathbb{Z}\rangle$. To define the target, consider $\Lambda$ as a free abelian group with basis $\{t^{r}:r\in\mathbb{Z}\}$, and take the quotient abelian group where $t^{r}$ and $t^{-r}$ are identified for all $r$. Taking the sum of contributions from all the double points $p$ gives
@@ 3. Turned twist-spun tori
In this section, we review Boyle’s construction [Boy93, Definition 5.1]. Let $K$ be a knot in $D^{3}$ such that $K$ intersects $S^{2}_{1/2}$ in $\{(0,0,1/2),(0,0,-1/2)\}$ transversely, and such that $K\setminus B^{3}_{1/2}$ is a boundary-parallel arc in $D^{3}\setminus B^{3}_{1/2}$ with endpoints on $S^{2}_{1/2}$. Furthermore, let $i$ and $j$ be integers. Let $R_{\theta}$ be rotation of $\mathbb{R}^{3}$ about the $z$-axis by angle $\theta$. Fix $\varepsilon\in(0,1/2)$. Choose an isotopy $\{\Phi_{t}:t\in I\}$ of $D^{3}$ such that $\Phi_{t}|_{S^{2}_{r}}=\text{Id}_{S^{2}_{r}}$ for all $r\in(1-\varepsilon,1]$, and
@@
for every $t\in I$. Write $\phi_{j}:=\Phi_{1}$. In addition, we define the diffeomorphism $\eta_{i}\colon S^{2}\times S^{1}\to S^{2}\times S^{1}$ via
@@
for each $\theta\in S^{1}$.
@@
where $S^{4}=(D^{3}\times\partial D^{2})\cup_{\eta_{i}}(\partial D^{3}\times D^{2})$ and $T_{K,i,j}:=K\times I$.
To see that the ambient manifold is indeed $S^{4}$, note that, if $i$ is odd, then it is obtained from
$S^{4}\cong\partial D^{5}\cong\partial(D^{3}\times D^{2})\cong(D^{3}\times\partial D^{2})\cup(\partial D^{3}\times D^{2})$
via a Gluck twist along the standard 2-sphere $S_{0}:=S^{2}\times\{0\}$. On the other hand, if $i$ is even, then $\eta_{i}$ is isotopic to $\text{Id}_{S^{2}\times S^{1}}$, and so we recover the standard $S^{4}$ directly. Intuitively, we can describe $T_{K,i,j}$ as follows. Let $K\subseteq\mathbb{R}^{3}_{>0}$ be a knot. Choose a 3-disc $D\subseteq\mathbb{R}^{3}_{>0}$ such that $\partial D$ intersects $K$ transversely in a pair of antipodal points and $K\setminus D$ is an unknotted arc. Then we obtain $T_{K,i,j}$ in $\mathbb{R}^{4}\subseteq S^{4}$ by rotating +$\mathbb{R}^{3}_{+}$ around $\mathbb{R}^{2}=\partial\mathbb{R}^{3}_{+}$, and while doing so, rotating $K\cap D$ around the line through $K\cap\partial D$ and simultaneously rotating $K$ around the $z$-axis. Boyle showed that the smooth isotopy class of $T_{K,i,j}$ only depends on $i$ through the parity of $i$. Indeed, the group of automorphisms of $S^{2}\times S^{1}$ modulo ones that extend to $S^{2}\times D^{2}$ is $\mathbb{Z}/2$, generated by $\eta_{1}$. Boyle calls $T_{K,0,i}$ the $j$-twist-spun torus of $K$ and $T_{K,1,j}$ the turned $j$-twist-spun torus of $K$. Boyle also showed that $\pi_{1}(S^{4}\setminus T_{K,i,j})\cong G/[G,\mu^{j}]$, where $G\cong\pi_{1}(S^{3}\setminus K)$ and $\mu\in G$ is the class of a meridian of $K$. Since $\mu$ normally generates $G$, we have $[G,\mu]\cong[G,G]$. This isomorphism can be seen using the commutator identities $[g,\mu^{x}]=\bigl{[}g^{x^{-1}},\mu\bigr{]}^{x}$ and $[g,\mu^{x}\mu^{y}]=[g,\mu^{x}][g,\mu^{y}]^{\mu^{x}}$, where superscripts indicate conjugation. Hence
@@ 4. Every turned 1-twist-spun torus is topologically unknotted
In this section, we study turned 1-twist-spun tori up to topological isotopy. Our main result is the following, restated from the introduction. See 1.1
@@
disjoint from $K$. This has self-intersection 00. Isotope the arc $K\setminus B^{3}_{1/2}$ relative to its boundary such that its interior only intersects the $z$-axis in $\{(0,0,3/4)\}$, where it is perpendicular to it. Let $\mu$ be a meridian of $K$ at $(0,0,3/4)$ of radius $1/8$. Then
@@
is a rim torus of $T_{K,1,1}$ over a longitude $\ell$ of $T_{K,1,1}$. Since $T_{\ell}$ is not a $\mathbb{Z}$-trivial surface, we surger it to a sphere. For this end, consider the arc
@@
is a disc in $S^{4}$ with boundary on $T_{\ell}$. Since $\eta_{1}$ applies a $2\pi$ rotation about the $z$-axis to $\mu$, a trivialisation of the normal bundle $\nu_{\partial D\subseteq T_{\ell}}$ does not extend to a section of the normal bundle $\nu_{D\subseteq E}$. The tangential framing of $\nu_{\partial D\subseteq T_{\ell}}$, transported to
@@
However, if we perform a boundary twist on $D$ along the curve $\partial D=\{(0,0,7/8)\}\times I$, then we can undo the twist. The resulting disc $D_{1}$ intersects $T_{\ell}$ once, but now a trivialisation of $\nu_{\partial D_{1}\subseteq T_{\ell}}$ does extend to $\nu_{D_{1}\subseteq E}$. If we surger $T_{\ell}$ along $D_{1}$, we obtain an immersed sphere $S_{1}$ with two double points, and we claim that $\mu([S_{1}])=\pm(1-t)\in\Lambda/\langle t^{r}\sim t^{-r}:r\in\mathbb{Z}\rangle$. We have that $D_{1}$ is embedded and framed, and intersects $T$ in one point, so, by part (i) of 2.5, we have that $\mu(S_{1})=\pm t^{k}(1-t)$ for some $k$. Since the double points arose from a boundary twist, it follows that one of the double points is local, i.e. the entire double point loop is contained in a copy of $D^{4}\subseteq E_{\Sigma}$. Hence, this double point loop is trivial. We deduce that $\mu(S_{1})=\pm(1-t)$ or ±$\pm(1-t^{-1})$. Given the indeterminacy in $\mu(S_{1})$, this completes the proof of the claim that $\mu(S_{1})=\pm(1-t)$.

### v3

@@ 1.2. Seifert surface union a ribbon disc
is surjective. In particular, every ribbon disc is homotopy ribbon; see Gordon [Gor81]. Let $K$ be a knot in $S^{3}$ with Seifert surface $F$ and homotopy ribbon disc $D$. Then $\Sigma:=F\cup D\subseteq S^{4}$ is a $\mathbb{Z}$-surface.
A slice derivative for a genus $g$ Seifert surface $F$ for a knot $K$ in $S^{3}$ is a collection of $g$ pairwise disjoint, homologically linearly independent simple closed curves that form a smoothly slice link and have self-linking zero [CHL10]. If $F$ has a slice derivative, we can perform surgery on $F$ in $D^{4}$ along a slice derivative and obtain a smooth slice disc $D$ for $K$. Then $F\cup D$ is smoothly unknotted; see Lemma 6.1. If $D$ is a ribbon disc that is not obtained from a Seifert surface $F$ by surgering along a slice derivative, Gabai has asked whether $F\cup D$ is topologically or smoothly unknotted.
Kauffman conjectured that every genus one Seifert surface for a slice knot admits a slice derivative. Cochran and Davis [CD15] have shown that there exists a ribbon knot admitting a genus one Seifert surface $F$ that has no slice derivative. Moreover, there is an example where $F$ is the unique minimal genus one Seifert surface, up to isotopy. The union of $F$ and a ribbon disc is therefore a potentially smoothly knotted $\mathbb{Z}$-torus in $S^{4}$. We show that it is topologically unknotted.
@@ 1.3. Tori with four critical points
Let $\Sigma\subseteq S^{4}$ be a smoothly embedded torus with four critical points. Suppose that the single-variable Alexander polynomial $\Delta_{L}$ of the 2-component link $L:=\Sigma\cap S^{3}$ is zero, where $S^{3}$ is the equator of $S^{4}$. Then $\Sigma$ is topologically unknotted.
@@
Let $\Sigma\subseteq S^{4}$ be a torus with four critical points, and suppose that the link $L:=\Sigma\cap S^{3}$ is split, where $S^{3}$ is the equator of $S^{4}$. Then $\Sigma$ is smoothly unknotted.
@@
Remark 1.6.
An interesting family of examples of $\mathbb{Z}$-surfaces that would be good candidates for potential future investigation are the following. Let $K$ be a knot in the 3-sphere with two non-isotopic genus one Seifert surfaces $F_{+}$ and $F_{-}$. Consider $S^{4}=S^{4}_{+}\cup_{S^{3}}S^{4}_{-}$, where $S^{4}_{\pm}$ is a copy of $D^{4}$, and push the interior of $F_{\pm}\subseteq S^{3}$ into $S^{4}_{\pm}$. The result $F_{+}\cup_{K}F_{-}$ is a genus two $\mathbb{Z}$-surface. Is it topologically, or even smoothly unknotted?
@@ Acknowledgements
We would like to thank Anthony Conway, Chris Davis, David Gabai, Jason Joseph, Maggie Miller, András Stipsicz, Zoltán Szabó, and Ian Zemke for helpful discussions and suggestions. We are also particularly grateful to Brendan Owens, who pointed out a mistake in our proof of an earlier version of 1.4, and to an anonymous referee for several very helpful suggestions on the exposition. MP was partially supported by EPSRC New Investigator grant EP/T028335/2 and EPSRC New Horizons grant EP/V04821X/2. For the purposes of open access, the authors have applied a CC-BY public copyright license to any Author Accepted Manuscript arising from this submission.
@@ 2. Standard intersection form implies unknotted
The elements of the standard basis $\{e_{i}\}_{i=1}^{2g}$ of $H_{2}(E_{\Sigma};\Lambda)\cong\Lambda^{2g}$ are represented by generically immersed spheres (i.e. locally a flat embedding and all self-intersections are transverse double points) in the smooth 4-manifold $\operatorname{Int}(E_{\Sigma})$ (it is smooth because it is an open subset of $S^{4}$), which we can see as follows. Apply the Hurewicz theorem to the universal cover of $E_{\Sigma}$ to see that $\pi_{2}(E_{\Sigma})\cong H_{2}(E_{\Sigma};\Lambda)$. Then apply the density of smooth immersions in $C^{0}(S^{2},\operatorname{Int}(E_{\Sigma}))$ [Hir76, Theorems 2.2.6 and 2.2.12], and finally the density of generic immersions in the space of smooth immersions [GG73, Chapter III, Corollary 3.3]. Since the $\{e_{i}\}_{i=1}^{2g}$ are represented by generically immersed spheres, so are $\{e_{i}\otimes 1\}_{i=1}^{2g}\subseteq H_{2}(E_{\Sigma};\Lambda)\otimes_{\Lambda}\mathbb{Z}\cong\Lambda^{2g}\otimes_{\Lambda}\mathbb{Z}$.
@@
for some $a_{ij}\in\mathbb{Z}$ such that the matrix $(a_{ij})_{i,j=1}^{2g}$ is invertible over $\mathbb{Z}$. Hence, the elements $\{e_{i}^{\prime}:=\sum_{j}a_{ij}e_{j}\}_{i=1}^{2g}$ generate $H_{2}(E_{\Sigma};\Lambda)$, because $(a_{ij})_{i,j=1}^{2g}$ is also invertible over $\Lambda$. Furthermore, the $e_{i}^{\prime}$ map to the rim tori. By taking $a_{ij}$ parallel copies of $e_{j}$ for $j\in\{1,\dots,2g\}$, linearly ordering all these spheres in an arbitrary way, and connecting each sphere with the next using a tube [FNOP24, Chapter 11] around an embedded framed arc that is coherent with the orientations of the spheres, we obtain a generically immersed sphere in $\operatorname{Int}(E_{\Sigma})$ representing $e_{i}^{\prime}$. ∎
@@
It will sometimes be helpful to compute the homological self-intersection numbers $\lambda([S_{i}],[S_{i}])\in\Lambda$ in Lemma 2.4 using the Wall self-intersection number $\mu(S_{i})$, so we recall this next; see [Wal99, Chapter 5]. Let $S\subseteq E_{\Sigma}$ be a $\mathbb{Z}$-trivial surface that is the image of a generic immersion into $E_{\Sigma}$. For each double point $p\in S$, we obtain a sign $\varepsilon_{p}$ and a double point loop $\gamma_{p}$ in $S$, with the latter well-defined up to orientation and homotopy in $E_{\Sigma}$; see e.g. [BKK+21, Section 11.3]. Thus each double point gives rise to a monomial ±$\pm t^{k}\in\Lambda/\langle t^{r}\sim t^{-r}:r\in\mathbb{Z}\rangle$. To define the target, consider $\Lambda$ as a free abelian group with basis $\{t^{r}:r\in\mathbb{Z}\}$, and take the quotient abelian group where $t^{r}$ and $t^{-r}$ are identified for all $r$. Taking the sum of contributions from all the double points $p$ gives
@@ 3. Turned twist-spun tori
In this section, we review Boyle’s construction [Boy93, Definition 5.1]. Let $K$ be a knot in $D^{3}$ such that $K$ intersects $S^{2}_{1/2}$ in $\{(0,0,1/2),(0,0,-1/2)\}$ transversely, and such that $K\setminus B^{3}_{1/2}$ is a boundary-parallel arc in $D^{3}\setminus B^{3}_{1/2}$ with endpoints on $S^{2}_{1/2}$. Furthermore, let $i$ and $j$ be integers.
Let $R_{\theta}$ be rotation of $\mathbb{R}^{3}$ about the $z$-axis by angle $\theta$. Fix $\varepsilon\in(0,1/2)$. Choose an isotopy $\{\Phi_{t}:t\in I\}$ of $D^{3}$ such that $\Phi_{t}|_{S^{2}_{r}}=\text{Id}_{S^{2}_{r}}$ for all $r\in(1-\varepsilon,1]$, and
@@
for every $t\in I$. Write $\phi_{j}:=\Phi_{1}$.
We define the diffeomorphism $\eta_{i}\colon S^{2}\times S^{1}\to S^{2}\times S^{1}$ via
@@
for each $\theta\in S^{1}$. Consider the decomposition
$S^{4}\cong\partial D^{5}\cong\partial(D^{3}\times D^{2})\cong(D^{3}\times\partial D^{2})\cup(\partial D^{3}\times D^{2}).$
Then the 4-manifold
$(D^{3}\times\partial D^{2})\cup_{\eta_{i}}(\partial D^{3}\times D^{2})$
is obtained from $S^{4}$ via a Gluck twist along the standard 2-sphere $S_{0}:=\partial D^{3}\times\{0\}$, and is hence diffeomorphic to $S^{4}$. On the other hand, if $i$ is even, then $\eta_{i}$ is isotopic to $\text{Id}_{S^{2}\times S^{1}}$, and so we again obtain the standard $S^{4}$.
@@
where $S^{4}\cong(D^{3}\times\partial D^{2})\cup_{\eta_{i}}(\partial D^{3}\times D^{2})$ and $T_{K,i,j}:=K\times I$.
Intuitively, we can describe $T_{K,i,j}$ as follows. Let $K\subseteq\mathbb{R}^{3}_{>0}$ be a knot. Choose a 3-disc $D\subseteq\mathbb{R}^{3}_{>0}$ such that $\partial D$ intersects $K$ transversely in a pair of antipodal points and $K\setminus D$ is an unknotted arc. Then we obtain $T_{K,i,j}$ in $\mathbb{R}^{4}\subseteq S^{4}$ by rotating +$\mathbb{R}^{3}_{+}$ around $\mathbb{R}^{2}=\partial\mathbb{R}^{3}_{+}$, and while doing so, rotating $K\cap D$ around the line through $K\cap\partial D$ and simultaneously rotating $K$ around the $z$-axis.
Boyle showed that the smooth isotopy class of $T_{K,i,j}$ only depends on $i$ through the parity of $i$. Indeed, the group of automorphisms of $S^{2}\times S^{1}$ modulo ones that extend to $S^{2}\times D^{2}$ is $\mathbb{Z}/2$, generated by $\eta_{1}$. Boyle calls $T_{K,0,i}$ the $j$-twist-spun torus of $K$ and $T_{K,1,j}$ the turned $j$-twist-spun torus of $K$.
Boyle also showed that $\pi_{1}(S^{4}\setminus T_{K,i,j})\cong G/[G,\mu^{j}]$, where $G\cong\pi_{1}(S^{3}\setminus K)$ and $\mu\in G$ is the class of a meridian of $K$. Since $\mu$ normally generates $G$, we have $[G,\mu]\cong[G,G]$. This isomorphism can be seen using the commutator identities $[g,\mu^{x}]=\bigl{[}g^{x^{-1}},\mu\bigr{]}^{x}$ and $[g,\mu^{x}\mu^{y}]=[g,\mu^{x}][g,\mu^{y}]^{\mu^{x}}$, where superscripts indicate conjugation. Hence
@@ 4. Every turned 1-twist-spun torus is topologically unknotted
In this section, we study turned 1-twist-spun tori up to topological isotopy. Our main result is the following, restated from the introduction.
See 1.1
@@
disjoint from $K$. This has self-intersection 00.
Isotope the arc $K\setminus B^{3}_{1/2}$ relative to its boundary such that its interior only intersects the $z$-axis in $\{(0,0,3/4)\}$, where it is perpendicular to it. Let $\mu$ be a meridian of $K$ at $(0,0,3/4)$ of radius $1/8$. Then
@@
is a rim torus of $T_{K,1,1}$ over a longitude $\ell$ of $T_{K,1,1}$.
Since $T_{\ell}$ is not a $\mathbb{Z}$-trivial surface, we surger it to a sphere. For this end, consider the arc
@@
is a disc in $S^{4}$ with boundary on $T_{\ell}$.
Since $\eta_{1}$ applies a $2\pi$ rotation about the $z$-axis to $\mu$, a trivialisation of the normal bundle $\nu_{\partial D\subseteq T_{\ell}}$ does not extend to a section of the normal bundle $\nu_{D\subseteq E}$. The tangential framing of $\nu_{\partial D\subseteq T_{\ell}}$, transported to
@@
However, if we perform a boundary twist on $D$ along the curve $\partial D=\{(0,0,7/8)\}\times I$, then we can undo the twist. The resulting disc $D_{1}$ intersects $T_{\ell}$ once, but now a trivialisation of $\nu_{\partial D_{1}\subseteq T_{\ell}}$ does extend to $\nu_{D_{1}\subseteq E}$.
If we surger $T_{\ell}$ along $D_{1}$, we obtain an immersed sphere $S_{1}$ with two double points, and we claim that $\mu([S_{1}])=\pm(1-t)\in\Lambda/\langle t^{r}\sim t^{-r}:r\in\mathbb{Z}\rangle$. We have that $D_{1}$ is embedded and framed, and intersects $T$ in one point, so, by part (i) of 2.5, we have that $\mu(S_{1})=\pm t^{k}(1-t)$ for some $k$. Since the double points arose from a boundary twist, it follows that one of the double points is local, i.e. the entire double point loop is contained in a copy of $D^{4}\subseteq E_{\Sigma}$. Hence, this double point loop is trivial. We deduce that $\mu(S_{1})=\pm(1-t)$ or ±$\pm(1-t^{-1})$. Given the indeterminacy in $\mu(S_{1})$, this completes the proof of the claim that $\mu(S_{1})=\pm(1-t)$.